​Claim: mu does not equal 7000​; alpha equals 0.03​; sigma equals 382. Sample​ statistics: x overbarequals 6800​, n equals 46 what is the critical value, or the critical value(s)?
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :   = 7000

Ha :    7000

= 6800

= 7000

= 382

n = 46

= 0.03

/ 2 = 0.03 / 2 = 0.015

Z/2 = Z0.015 = 2.17

Critical value : -2.17 and +2.17

P-value <

Reject the null hypothesis .